The sum of four consecutive integers is 34. What is the smallest of the four integers?
Answer by macston(5194)

n=smallest integer; n+1=second integer; n+2=third integer; n+3=fourth integer
n+(n+1)+(n+2)+(n+3)=34
4n+6=34
4n=28
n=7 ANSWER: The smallest integer is 7
CHECK:
7+8+9+10=34
34=34